# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v1

@@ 6.1 Systems for RAG
Several prior works have attempted to build systems for RAG. RAGCache [44] proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token. However, this approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency [4]. Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores. Similarly, TurboRAG [60] precomputes KV caches from the data store, but it also only optimizes prefill latency.
@@
RaLMSpec [98] proposes speculative retrieval and batched verification, Chameleon [41] proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG [42] proposes an algorithm-system co-design technique to overlap the retrieval and generation process by modifying the algorithms. However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and are not applicable to modular RAG applications that are widely used now, as discussed in §2.1.
@@
Speculative RAG [87] introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency. APIServe (InferCept) [3] proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval. However, this work again does not focus on optimizing retrieval latency.
@@
6.2 Systems for Compound LLM Applications
Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities [95, 14, 85]. LLMCompiler [52] is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling. AI Metropolis [90] accelerates LLM-based multi-agent simulations with out-of-order execution. ALTO [76] improves the performance of complex AI systems by streaming partial results. Some open-source libraries provide graph-based representation for programming this kind of application [22, 56]. RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
6.3 Vector Index
Vector index is gaining growing attention as an essential component of RAG [16, 29], and many works have been proposed to improve the efficiency with system or algorithm techniques. ScaNN proposes an anisotropic quantization method for better efficiency [32]. DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity [19, 38]. Other prior works propose hardware acceleration of vector index with GPUs [46], FPGAs [96, 40], TPU [20], or ray tracing hardware [104, 59]. BANG proposes a method to scale graph-based ANN beyond GPU memory [49], and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining [99]. However, these methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth. Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
7 Conclusion

### v2

@@ 6.1. Systems for RAG
RAGCache (Jin et al., 2024a) proposes a caching system that stores KV caches from datastore chunks in an order-aware manner to improve the time to the first token. This approach only reduces prefill latency, leaving retrieval and decode times unaffected, despite the fact they typically dominate total latency (Agrawal et al., 2024). Moreover, it assumes repeated use of the same document across multiple requests, limiting its scalability over large data stores. Similarly, TurboRAG (Lu et al., 2024) precomputes KV caches from the data store, but it also only optimizes prefill latency. CacheBlend introduces a selective KV-cache fusion technique for RAG to reuse pre-computed caches (Yao et al., 2024).
@@
RaLMSpec (Zhang et al., 2024b) proposes speculative retrieval and batched verification, Chameleon (Jiang et al., 2023d) proposes a CPU-GPU-FPGA heterogeneous architecture for accelerating the retrieval process, and PipeRAG (Jiang et al., 2024) is an algorithm-system co-design technique to overlap the retrieval and generation by modifying the algorithms. However, these works focus on the paradigm of RAG that retrieves documents once every few tokens and do not apply to modular RAG applications that are widely used now, as discussed in §2.1.
@@
Speculative RAG (Wang et al., 2024b) introduces drafting by smaller LLMs to reduce RAG latency, but it does not target the retrieval latency. APIServe (InferCept) (Abhyankar et al., 2024) proposes a novel KV cache management strategy that can support the interception of LLM generation by other workloads, including retrieval. However, this work again does not focus on optimizing retrieval latency. EdgeRAG (Seemakhupt et al., 2024) reduces the memory requirement of retrieval by re-generating and caching document embeddings at runtime, targeting extremely resource-constrained environments. RAGO (Jiang et al., 2025) introduces an abstraction for RAG to automatically pick task placement, resource allocation, and batching policies.
@@
6.2. Systems for Compound LLM Applications
Apart from RAG, there is a growing interest in compound or agentic LLM applications, where multiple LLM calls and other applications are combined to serve complex functionalities (Zaharia et al., 2024; Berger and Zorn, 2024; Wang et al., 2024a). LLMCompiler (Kim et al., 2023a) is a framework that optimizes the execution of multiple functions in large language models by enabling parallel function calling. AI Metropolis (Xie et al., 2024) accelerates LLM-based multi-agent simulations with out-of-order execution. RAG is a specific type of application in this broader direction, and we propose systems techniques to optimize its execution latency, focusing on the characteristics of retrieval workload.
6.3. Vector Index
Vector index is a key component of RAG (Briggs et al., [n. d.]; Gao et al., 2023b), and many works have been proposed to improve their efficiency. ScaNN proposes an anisotropic quantization method for better efficiency (Guo et al., 2020). DiskANN and SPANN propose memory-disk hybrid indexing systems that work beyond the limitation of memory capacity (Chen et al., 2021; Jayaram Subramanya et al., 2019). Other prior works propose hardware acceleration of vector index with GPUs (Johnson et al., 2019), FPGAs (Zeng et al., 2022; Jiang et al., 2023a), TPU (Chern et al., 2022), or ray tracing hardware (Zhu, 2022; Liu et al., 2023). BANG proposes a method to scale graph-based ANN beyond GPU memory (Khan et al., 2024), and Rummy allows the index to scale beyond GPU memory capacity with reordered pipelining (Zhang et al., 2024a). These methods either require algorithm modifications or are bottlenecked by CPU-GPU bandwidth. Our proposal, focuses on the context of modular RAG applications, where queries for retrieval are usually generated by LLMs, and optimizes the latency and the GPU memory consumption without altering the algorithm of the IVF index.
7. Conclusion